CHARGE REDISTRIBUTION IN THE β-NAPHTHOL-WATER COMPLEX AS MEASURED BY HIGH RES-OLUTION STARK SPECTROSCOPY IN THE GAS PHASE.^{a}

\underline{ADAM J. FLEISHER}, DAVID W. PRATT, Department of Chemistry, University of Pittsburgh, Pittsburgh, PA 15260; ALESSANDRO CEMBRAN and JIALI GAO, Department of Chemistry, University of Minnesota, Minneapolis, MN 55455.

The extensively studied photoacid β-naphthol exhibits a large decrease in pK_{a} upon irradiation with ultraviolet light, in the condensed phase.^{b} β-naphthol is almost 10 million times more acidic in the excited electronic state, compared to the ground state. Motivated by this fact, we report here the measurement of the electronic dipole moments of the β-naphthol-water complex in both electronic states, from which estimates of the charge transfer from solute to solvent in both states will be made. Comparisons to ab initio and density functional theory calculations will also be reported.^{c}

aWork supported by NSF (CHE-0911117).

bN. Mataga and T. Kubota, Molecular Interactions and Electronic Spectra (Marcel Dekker, New York, 1970).

cY. Mo, J. Gao, S.D. Peyerimhoff, J. Chem. P hys. 112, 5530 (2000).